In the last week we have had a few updates for two of our instruments. For VIRUS we are up to 29 working spectrographs. We actually have several more units but some of the oldest and slightly mis-behaving units have been sent back to Austin for realignment and recommissioning. For HPF we are thrilled to announce that a NIST laser comb has been installed in the calibration room in the HET basement. This allows us to send a picket fence of spectral features through a separate fiber next to the science fibers. During the data reduction and analysis of the HPF spectra they can look at the position of these pickets and determine how the instrument might be subtly moving and correct for it. So far it seems like it is working great.